### v1

# Improving Zero-noise Extrapolation for Quantum-gate Error Mitigation using a Noise-aware Folding Method

Abstract
The current thousand-qubit processors mark a substantial advance in hardware. Yet, hardware limitations prevent quantum error correction (QEC), necessitating reliance on quantum error mitigation (QEM). Our paper presents a noise-aware folding method that improves Zero-Noise Extrapolation (ZNE) by estimating noiseless values from noisy results. Unlike traditional ZNE methods, which assume a uniform error distribution, our method redistributes the noise using calibration data based on hardware noise models. By employing noise-adaptive compilation and optimizing the qubit mappings, our approach enhances the ZNE accuracy of various quantum computing models. Recalibrating the noise amplification to address the inherent error variations, promises higher precision and reliability in quantum computations. This paper highlights the uniqueness of our method, summarizes noise accumulation, presents the scaling algorithm, and compares the reliability of our method with those of existing models using linear fit extrapolation. Relative to the existing folding methods, our method achieved a 35% improvement on quantum computer simulators and a 26% improvement on real quantum computers compared to existing folding methods, demonstrating the effectiveness of our proposed approach.
preprint: APS/123-QED
I Introduction
Recently, quantum computing (QC) technology has entered the era of noisy-intermediate scale quantum computing (NISQ), highlighted by IBM’s recent release of processors housing over a thousand qubits [1, 2]; meanwhile, robust crosstalk chips using tunable-coupler technology, can operate up to 133 qubits [3, 4, 5]. Despite these advancements, the capacity of current hardware is inadequate for implementing quantum error correction codes but is projected to become adequate by the 2030s [6, 7]. In the meantime, QC systems employ QEM methods, often requiring additional quantum and classical resources as a trade-off to enhance QC output fidelity.
Numerous QEM techniques for alleviating diverse sources of error in QC have been proposed [8]: probabilistic error cancelation for mitigating decoherence  [9, 10], ZNE for imperfect gates, techniques for mitigating measurement errors  [11, 12, 13], dynamical decoupling [14, 15], quantum optimal control [16, 17], randomized compiling [18, 19], Pauli-frame randomization [20, 21], and other techniques [22, 23, 24, 25, 21, 26]. The present study focuses on the well-established ZNE technique.
ZNE was concurrently introduced and extensively demonstrated in various applications involving systems of up to 127 qubits [27, 28]. In ZNE, a quantum program is amplified to different noise levels through gate or pulse-level stretching, which intentionally increases the program’s noise. Subsequently, the amplified results are extrapolated to estimate the noiseless values. Formally, the program is amplified by multiple scale factors $\lambda$. When $\lambda=1$, the program operates at its original error rate. When $\lambda$ $>0$, an additional error, such as an identity unitary combination of the original gate, is introduced. The results obtained at different noise levels of $\lambda$ are collected and extrapolated to $\lambda$ $=0$, effectively eliminating the noise. Therefore, $\lambda$ represents the noise level affecting any physical quantity during quantum computation.
Researchers have proposed various methods for amplifying noise in quantum programs. For instance, the approach described in [29] stretches gate durations to the desired levels using pulse-level control. Under ideal conditions, this stretching does not alter the quantum system’s state, but under noisy conditions, these modifications are classified as errors. Moreover, as this method requires a high degree of abstract control and calibration at the pulse level in quantum computers, it is not easily implementable across most of the existing QC systems. In constrast, unitary folding [30] amplifies the noise using gate-level control, which is available in all gate-based computation model QCs. This technique builds upon a simple concept: replacing the unitary operation $U\rightarrow U\left(U^{\dagger}U\right)$, where $\left(U^{\dagger}U\right)$ is an identity. In an ideal scenario, this replacement remains a $U$ operation, but errors corresponding to U in the quantum system amplify the errors in the $U$ operation within the modified circuit. Giurgica-Tiron et al.  [30] proposed unitary folding with a fold from the left (which folds each unitary gate independently) and at random folding methods (which randomly selects a subset of individual gates randomly as a block of unitary gates and replicates the entire block with $U\left(U^{\dagger}U\right)$. This method requires no knowledge of the QC underlying noise model. Although these scaling methods effectively amplify noise in the original input quantum program, they neglect the sources of error imbalance within the QC system, potentially leading to biased results and errors that may compromise the accuracy of the extrapolation results.
To resolve this problem, we introduce a noise-aware folding method that adjusts the quantum circuit, aiming to balance the error rate across all logical qubits by leveraging the calibration data noise model, which is freely available on select QC systems. This approach comprehensively analyzes the calibration data noise model, thereby redistributing the noise amplification process across the quantum circuit. Through an algorithmic design, our approach strategically adjusts the scaling factor for each gate operation, balancing the distribution of gate error rates across all logical qubits. Addressing the error source imbalances, mitigates bias-induced errors, thereby enhancing the accuracy of ZNE extrapolated results from the quantum program.
Moreover, the approach is adaptable and therefore versatile across various quantum computing models. Seamlessly integrated with gate-based computation models, it comprehensively caters to the inherent complexities of contemporary quantum hardware. The integration of a noise-aware folding method enhances the precision and reliability of quantum computations. The amplification process can be recalibrated to accommodate the intricate error landscapes within quantum systems, promising more accurate and dependable quantum programming, and bridging the gap between theoretical expectations and practical outcomes. The contributions of our study are highlighted below:
  • Our noise-aware folding method is tailored to redistribute noise across quantum circuits, thereby enhancing the accuracy of ZNE by addressing the inherent error variations within quantum systems.
  • We seamlessly integrate our approach with gate-based computation models and optimize the qubit mappings using hardware noise models, providing versatility mitigating bias-induced errors, and enhancing the reliability of quantum computations.
  • Our novel algorithm dynamically adjusts the error rates within the quantum circuit, leveraging calibration data and noise-adaptive compilation methods to ensure a balanced and controlled scaling process.
  • We conduct rigorous experiments across various full noise models and real quantum computers, showcasing the method’s consistent performance in simulations, highlighting the challenges of scaling to larger circuits, and emphasizing the disparity between simulated and real quantum computer executions.
Figure 1: Controlled-not (CNOT) operation as identity insertion of unitary U1 and U2 operations using two qubits.
The remainder of this is organized as follows. Section II.1 briefly discusses the background and related QEM methods using ZNE techniques, which motivated our research. Section III explains the noise-aware folding method, along with its compilation and execution scheme. Section IV analyzes the performance results of our approach and previous methods. Our noise-aware folding method is further discussed in Section V and conclusions are presented in Section VI.
II Background and Related Works
This section briefly provides the fundamentals of our proposed method. We discuss the quantum error mitigation (QEM), noise scaling methods, and noise extrapolation models.
II.1 Quantum Error Mitigation
Quantum computers are susceptible to various sources of noise and imperfections including crosstalk, which can affect the accuracy, reliability, and fidelity of QC results. Utilizing a set of techniques and strategies to minimize quantum computational errors, QEM estimates the true expectation value $\langle A\rangle$ from the approximated value $A$ impacted by errors $\psi$.
$\langle A\rangle=\operatorname{Tr}(A\psi)$ (1)
QEM strategically minimizes the impact of errors at various stages of quantum processes such as ZNE. Ultimately, it aims to optimize the performance of quantum algorithms and facilitate their seamless integration into practical QC environments. As one of the best-estimated QC techniques, ZNE is used for understanding the behavior of quantum algorithms in the absence of noise. In ZNE, quantum computations are run at different noise levels and the outcomes are extrapolated to a hypothetical noiseless scenario (the zero-noise limit). By analyzing the results obtained at various noise levels, we can predict the performance of a quantum algorithm without noise interference. This extrapolation provides insights into the intrinsic properties and capabilities of the quantum algorithm, clarifying its idealized behavior. Essentially, ZNE can separate the impact of noise from the true potential of a quantum algorithm.
II.2 Noise Scaling Methods
ZNE methods measure a given observable at varying levels of noise. Using the measured dependence on the noise, they then extrapolate the result to the expected noiseless value. Instruction-level noise in current digital quantum computers dominantly arises from the two-qubit entangling controlled-NOT (CNOT) gate, and the dominant noise channel is the two-qubit depolarizing channel [31]. In a two-qubit scenario, the depolarizing channel is given by a quantum operation acting on the system’s density matrix $\rho$:
$\mathcal{E}(\rho)=(1-\epsilon)\rho+\frac{\epsilon}{4}I$ (2)
where $\epsilon$ is the error rate parameter and $I$ is the 2 × 2 identity matrix.
We denote the density matrix of a single noisy (depolarizing) CNOT gate on two-qubit $i$ and $j$ by $\rho_{ij}$ and the unitary operator corresponding to the CNOT gate by $U_{c}$. As the action of two CNOT gates is equivalent to the identity operation (see Figure 1 an odd scaling factor $\lambda$ of CNOT gates must be added to the same qubits as follows
$\operatorname{CNOT}_{ij}^{\lambda}[\rho]=(1-\lambda\epsilon)U_{c}\rho U_{c}+\frac{\lambda\epsilon}{4}\rho_{ij}\otimes I_{ij}+O\left(\epsilon^{2}\right)$ (3)
Unitary folding modifies the quantum circuit by incorporating additional operations or ’folds’ that counteract the effects of noise. This strategically introduces corrective operations that enhance the fault-tolerance of the quantum computation, ultimately mitigating the impact of errors.
Parameterized noise scaling counteracts the effects of noise by adjusting the parameters of quantum gates in a controlled manner to account for and counteract the effects of noise. By dynamically scaling the parameters of quantum gates based on the noise characteristics, this technique optimizes the overall performance of the quantum algorithm in noisy environments.
Several methods for scaling quantum circuits have been reported [30, 29, 32, 33]. The two well-known approaches of Giurgica-Tiron et al. [30], i.e., fold from the left and random folding, manipulate noise solely at the gate 0 level to enable ZNE.
  • Noise Scaling from the Left: This approach selects a specific subset of individual gates or layers in a quantum circuit from the left-hand and subject it to noise scaling operations, starting from the leftmost side of the circuit until the end of the quantum circuit. This approach enables targeted noise scaling of specific portions of the circuit, potentially controlling the overall noise characteristics and error mitigation in a controlled manner.
  • Noise Scaling at Random: This approach randomly selects a subset of individual gates or layers for noise scaling operations. Random gate selection enables uniform sampling of the input circuit, smoothing the convergence to a specific scaling factor. By introducing randomness into the gate selection process, this approach can achieve effective noise scaling and error mitigation in QC.
II.3 Noise Extrapolation Models
Extrapolation models for ZNE in quantum computing have been comprehensively discussed elsewhere [22, 30, 10]. The choice of the model depends on the strength of the noise and the number of gates in the circuit. The four different extrapolation models are summarized below.
  • The linear model [22] assumes that the expectation value of a quantum circuit scales linearly with the number of gates. The linear model is simple and can be useful for estimating the zero-noise limit expected value of the circuit when the noise is weak.
  • The exponential model [22] assumes that the noise in a circuit scales exponentially with the number of gates. This model can be used with unitary folding for exponential scaling of the depolarizing parameters of each gate. The obtained parameters are then used for fitting and extrapolating the expectation value of the circuit.
  • The adaptive-exponential model  [30] an extension of the exponential extrapolation model, allows the scaling factor to vary with the number of gates. Estimating the scaling factor using a Bayesian approach, this method improves the accuracy of the extrapolation and reduces the variance of the estimator.
  • The Richardson model [10] maximizes the order of the polynomial extrapolation to the maximum given the number of data points. When the noise is strong and the polynomial order is high, this model improves the accuracy of the extrapolation and reduces the variance of the estimator.
However, unitary folding [30] might not effectively amplify systematic nonuniform noise on the target quantum hardware. Giurgica-Tiron et al. demonstrated that the adaptive-exponential extrapolation method maximizes the accuracy and minimizes the variance, although the exponential, linear, and Richardson extrapolations are useful in different scenarios.
Figure 2: Overall process of using the noise-aware folding. A traditional qubit mapping process is applied to get a mapped circuit. Using the mapped circuit with the target hardware error rate, we can accumulate the noise on the circuit. Finally, we apply the noise-aware folding method to scale the quantum circuit.
III Proposed Noise-aware Folding ZNE
III.1 Overview
Previous methodologies adjusted the gate count of the quantum circuit using the scale factor ($\lambda$), employing techniques such as gate-level unitary folding [30], identity gate insertion [32], and pulse stretching [29] to account for noise scaling. Although these approaches require no specific hardware noise models, they assume uniform noise scaling across the circuit, ignoring the variations in error levels among qubits within the quantum system. In real systems, where the error rates differ at distinct gate levels, these scaling methods cannot obtain uniform scaling factors. Such nonuniform scaling factors can impede the convergence of the extrapolation model toward a zero-noise state effectively.
To tackle these problems, we propose leveraging the noise model of the target hardware to scale the noise on the quantum circuit for ZNE. Our approach regards $\lambda$ not solely as a gate-count scaling factor, but as an error-rate modifier that directly affects the quantum circuit. Quantum computers demand periodic calibration, meaning that calibration data are available for error mitigating by methods such as noise-adaptive compilation [34]. Our approach optimizes the mapping using the noise-adaptive compilation proposed by Murali et al. [34], which favors qubits with high resilience. This method maps logical qubits to physical qubits with low noise levels and minimal distance, thereby reducing the SWAP gate requirements. Employing our noise-aware folding algorithm, we then scale the approximate error rates within the transpiled circuit effectively. The following section explains the specifics of our noise-aware folding method.
III.2 Noise Accumulation
Figure 2 illustrates a quantum circuit and its associated target hardware data, which are input to our proposed method. First, we transpile the quantum circuit using the noise-adaptive compilation mentioned in Section III, efficiently mapping the quantum circuit onto the target hardware by selecting the most robust qubits with low error rates. Although noise is intentionally added to the circuit, noise-adaptive mapping is crucial for ensuring reliable operations, minimizing qubit movement, and reducing the need for SWAP gates, which largely affect the performances of folding methods. To leverage this advantage, a deeper understanding of quantum computer noise—often overlooked in previous ZNE research—is crucial.
The transpiled circuit accumulates noises into an $n$$\times$ $n$ matrix denoted as $qc\_matrix$ in Algorithm 1, where $n$ represents the number of qubits. Along each off-diagonal (where $i$ and $j$ are the rows and column indices, respectively), we aggregate the noise for two-qubit gates between each qubit pair (where $i$ and $j$ denote the indices of the control target qubit, respectively, with $i\neq j$). The noise of the one-qubit gates is accumulated along the diagonal matrix ($i=j$). Because the error rate of a two-qubit gate is independent of the control or target qubit, the noise in the two-qubit gates is accumulated only between qubit pairs located on the upper diagonal of the matrix.
The error rate for each pair of qubits $i$ and $j$, the error rate is aggregated only if these qubits are physically connected and if two-qubit gates exist between them. Once the error rates of the quantum circuit are aggregated into the error rate matrix, the circuit can be scaled at each scale factor.
Input:
$circuit$: input transpiled circuit,
$scale\_factors$: a list of scale factors $\lambda$,
$backend$: target backend quantum computer
Output:
$folded\_circuits$ : a list of scaled circuits
1 Function  fold_noise_aware
2       $cx\_error\_dict$leftarrow GetErrorRateFrom($backend$);
3       $qc\_matrix$leftarrow AccumulateCircuitErrorRate($circuit$);
4       $folded\_circuit$leftarrow empty list;
       // scale the circuit in every $\lambda$ scale factor
5       for $scale$ in $scale\_factors$ do
6             $scaled\_matrix$leftarrow $qc\_matrix$ $scale$;
7             $qc\_rate$leftarrow GetHighestRateIn($qc\_matrix$);
8             $scaled\_rate$leftarrow GetHighestRateIn($scaled\_matrix$) $adjust\_rate$leftarrow ($qc\_rate$ + $scaled\_rate)$ /// $2$;
             // Scale each pair of the qubit on the upper diagonal of the $qc\_matrix$
9             for $i\leftarrow 0$ to $circuit.num\_qubits$ do
10                   for $j\leftarrow$ $i$ to $circuit.num\_qubits$ do
11                         $cur\_rate$leftarrow $qc\_matrix$[i][j];
                         // Perform unitary folding until the adjust rate is reached
12                         while $cur\_rate<adjust\_rate$ do
                               // Get the qubit error rate using control (i) and target (j) qubit as index.
13                               $gate\_rate$leftarrow $cx\_error\_dic[i,j]$;
                               // For every unitary folding insertion, twice the gate $(U)$ error rate is added for $U^{\dagger}U$
14                               $cur\_rate$ += $gate\_rate*2$;
                               // Apply CNOT gate folding to pair i,j
15                               $circuit.cx(i,j)$;
16                               $circuit.cx(i,j)$;
17
18                         end while
19
20                   end for
21
22             end for
23            $folded\_circuit$.insert($circuit$)
24       end for
25      return $folded\_circuit$;
26 end
ALGORITHM 1 Noise-aware Folding
III.3 Noise-aware Folding
As previously discussed, $\lambda$ in our methodology is leveraged to amplify the noise accumulation and is not directly correlated with the circuit’s gate count. This approach involves a nuanced computation (Algorithm 1). By integrating the quantum circuit error rate matrix ($qc\_matrix$ on Line 3) with a scaling constant, we derive the $scaled\_matrix$ on Line 6, effectively delineating the desired error rate augmentation of the scaled $\lambda$ circuit.
To control the pace of noise amplification, our strategy also calculates the average disparity between the maximum error rates within the $scaled\_matrix$ and the ongoing $qc\_matrix$ (Lines 6-9 of Algorithm 1). The computed balanced error rate termed the $adjust\_rate$ on Line 9, is a regulatory mechanism that caps the rate at which errors accumulate.
Next, the error rate is amplified through the unitary folding technique, which substitutes $U$ with $U(U^{\dagger}U)$ to introduce supplementary gates. For every qubit pair within the $qc\_matrix$, this procedure iterates through the gate increments until the existing error rate of the qubit pair ($cur\_rate$) approaches or slightly surpasses the designated $adjust\_rate$ (Lines 10-17 of Algorithm 1).
The method concludes after evaluating all pairs along the upper diagonal matrix. The error rate in resultant circuit is approximately equivalent to the predefined scale $\lambda_{n}=(\lambda_{1}+(\lambda_{1}*\lambda_{n}))/2$.
Importantly, because error rates are available only for gate operations in the noise model, this study exclusively accumulates the noise in two-qubit gates. Nevertheless, our proposed methodology potentially enables seamless extension to all types of gates, provided that adequate error rate calibration data exist within the noise model.
IV Evaluation
This section compares the results of our noise-aware folding method with those of other extrapolation models, including the existing method referenced in [30], using the linear fit extrapolation model.
Figure 3: The circuit configuration designed to benchmark ZNE using various folding methods begins with initializing each qubit to the state $|0\rangle$. Subsequently, a NOT (X) gate is applied to the first qubit, followed by successive CNOT gates between adjacent qubits (Qubit 1 and 2, 2 and 3, …, up to n-1 and n). This sequence results in all qubits being measured in the state $|1\rangle$.
IV.1 Experimental Setup
(a) ibm_mumbai noise model result.
(b) ibm_cairo noise model result.
Figure 4: Comparison of expectation values (vertical axis) obtained through using ZNE linear fit model with different folding methods on a quantum computer simulator, each with its respective noise model. (a) and (b) display results utilizing noise models from ibm_mumbai and ibm_cairo, respectively. The evaluation involves three folding methods: fold from left, fold at random, and noise-aware folding applied to the circuit depicted in Figure 3, spanning from 2 to 27 qubits (horizontal axis). The scaling factor $\lambda$ of $[1,1.5,2,2.5]$ was utilized for each folding method at every qubit count. The ‘Exact Simulation’ represents results without any applied noise model, while the ‘Unmitigated’ results showcase the circuit execution using a noise model without error mitigation. Each result is represented by a distinct symbol and color, as shown in (a).
Our proposed method was validated in various full-noise models from quantum computers available on Qiskit [35]. In these evaluations, we evaluated both the setup circuit in Figure 4 and the Bernstein-Vazirani (BV) circuit as a benchmark [36]. We chose this circuit for benchmarking our proposed method for specific reasons. At a logical level, it presents a relatively straightforward circuit that alters the states of individual qubits. When errors or noise are introduced to this circuit, they are nonnegligibly propagated between qubits and become significant due to their correlation and the necessity to change subsequent qubit states. Any error during execution notably impacts the subsequent qubit within the circuit, depending on the point of its incidence. Moreover, owing to the connectivity constraints, this circuit is not easily executable on quantum computers. Accommodating these limitations in quantum computer connectivity becomes increasingly intricate with the increasing number of qubits.
For each noise model, we ran the quantum circuit with different scaling parameters for comparison purposes. Along with our noise-aware folding method, we employed both from-left and at-random folding methods in [30] implemented on the Mitiq framework [37]. However. owing to the nature of ZNE, we must preserve the folded gates in the folded circuits, which precludes optimization such as gate cancelations before executing the quantum circuit at the compilation level. Given the complexity of our benchmark circuit (Figure 3) and the connectivity limitations of the quantum computer, we must introduce numerous SWAP gates as the number of qubits increases. To avoid this problem and ensure a fair comparison, we preemptively transpile the quantum circuit using Qiskit at the highest $optimization\_level=3$ before initiating the folding process in each method. The resulting transpiled circuit was input to each folding method, each with its own processes for extrapolation.
In addition, all experimental circuits were executed on real quantum computers (see Subsection IV.3). All comparison methods, on both the simulators and real quantum computers, employed the same parameters and compilation process. The subsequent subsection delves into the evaluation results.
IV.2 Result on Simulators
Our experiments were conducted on the currently available 27-qubit system, making use of the full-noise models implemented in Qiskit. Each experiment was conducted five times with different qubit counts and folding methods. Reported are the average outcomes of the five experiments for each folding method. The following graphs present the result of the linear fit extrapolation model, which demonstrated notably reliable results across all qubit counts.
Figure 4 depicts the outcomes of executing the benchmark circuit Figure 3 on three full-noise models, namely, ibm_mumbai and ibm_cairo, using a simulator. As shown in panels (a) and (b) of this figure, the expectation decreased with increasing qubit counts. Applying the unmitigated, fold from left, random folding, and noise-aware folding to the benchmark circuit in the ibm_mumbai noise model, the expectation values respectively decreased from 0.943, 0.952, 0.949, and 0.959 in the 2-qubit circuits to 0.275, 0.347, 0.350, and 0.397 in the 27-qubit circuit (Figure 3(a)).
The expectation values of all execution methods, (including the unmitigated method) were high in the ibm_cairo noise model than in the ibm_mumbai (c.f. Figure 3(a) and  3(b)). This discrepancy stems from differences in the noise model parameters. Specifically, ibm_mumbai yields a significantly higher error rate than the other models.
Figure 5 presents the experimental results of a 14-qubit BV circuit using the full-noise model simulation from ibm_cairo. The expectation value reduced with increasing scaling factor $\lambda$ in our proposed method, but showed no consistent drop in the left and random folding methods. This result demonstrates that in our proposed method, the extrapolation model accurately extrapolates to the zero-noise value.
IV.3 Result on Real Quantum Computers
Figure 6 presents the expectation values on the benchmark circuit in Figure 4, obtained through experiments on a real quantum computer ibm_algiers. The results, which mirror those of Figure 4, were averaged over five executions of each folding method. Notably, the expectation values of the unmitigated, fold from left, fold at random, and noise-aware folding methods declined from 0.979, 0.98, 0.99, and 0.979, respectively, in the 2-qubit circuit to 0.005, 0.021, 0.011, and 0.25, respectively in the 24-qubit circuit. The proposed method consistently outperformed the existing methods from the 2 to 16 qubits and performed comparably to the existing methods beyond the 16-qubit threshold. This discrepancy can be explained by the noticeably faster decline of the expectation value on scaled circuits in our proposed method than in the existing ones. At higher $\lambda$ values, certain folded circuits reach a 0 expectation value, rendering the extrapolated results unreliable.
Comparing executions on real quantum computers to simulations, the overall results deteriorate significantly with increasing qubit counts on real quantum computers. This disparity arises because the noise model does not encompass all noise sources present in real quantum computers during simulations. Consequently, the results are less reliable on real quantum computers.
When the BV circuit was executed on a real quantum computer, the success rate was consistently below 1%. This low success rate translates to an expectation value of less than 0.01 for the unmitigated circuit. Consequently, in folded circuits where the noise is increased, the success rate of the BV circuit approaches 0% and the results become meaningless on this circuit.
Figure 5: Comparison of the expectation value using different folding methods on a Bernstein-Vazirani (BV) [36] circuit with 14-qubit on ibm_cairo noise model simulation. The scale factor shows the drop in expectation value on different folding methods as the circuit gets nosier while $\lambda=0$ is the extrapolated expectation value where the zero-noise is assumed.
Figure 6: Expectation value results on a real quantum computer ibm_algiers with the exact same setup and scaling factor configuration from the result in Figure 4.
V Discussion
We applied other extrapolation methods, namely, exponential fit [22], adaptive-exponential fit from [30], and Richardson’s method [10], to ZNE execution on the above-mentioned circuit folding methods. However, our benchmarking results demonstrated no consistent substantial improvements or more reliable outcomes over those of linear fitting. Indeed, the models were often degraded by extreme overfitting or underfitting, leading to nonconvergence in some instances. Although improvements from those of linear fit extrapolation were sporadically observed, the results of these alternative fittings were too inconsistent for practical execution, particularly in scenarios involving variational-based quantum algorithms.
Notably, the exponential fit and adaptive-exponential fit methods produced identical outputs across all folding methods, including our proposed approach. Therefore, the efficacy and discriminative power of these extrapolations in distinguishing the performances of different extrapolation techniques is questionable. To evaluate our noise-aware folding method against the existing approaches, we conducted comprehensive experiments using linear fit extrapolation. Simulations on various full-noise models demonstrated that our proposed method consistently and efficiently performs across different quantum circuit complexities. Notably, it maintained high extrapolation accuracy as the circuit complexity increased.
Real quantum computer executions also delivered promising results, particularly on smaller circuits. However, at higher qubit counts, the expectation values fell more rapidly with circuit scaling than in the simulation results, affecting the reliability of the extrapolated results. The disparity between the simulation and real quantum computer executions highlights the challenges posed by the noise models, which do not capture all noise sources in actual hardware.
VI Conclusion
Our noise-aware folding method more effectively mitigates quantum errors during computations than the existing methods. By dynamically adjusting the error rates based on hardware noise models, it promises to improve the precision and reliability of quantum computations. The fidelity improvements over the compared methods reached 35% and 26% on quantum computer simulators and real quantum computers, respectively. Although demonstrating robustness in simulations and smaller-scale executions on real quantum computers, the proposed method was challenged by scaling to larger circuits, necessitating further exploration and optimization to broaden its applicability. Owing to its adaptability and versatility, the proposed method can potentially bridge the gap between theoretical expectations and practical outcomes in QC. Further refinements and adaptations are imperative to enhance scalability and reliability, propelling quantum error mitigation methods towards realizing error-free quantum computations.
Acknowledgements.
This research was partly supported by Quantum Computing based on Quantum Advantage challenge research (RS-2023-00257994) through the National Research Foundation of Korea(NRF) funded by the Korean government (MSIT) and Institute for Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. 2020-0-00014, A Technology Development of Quantum OS for Fault-tolerant Logical Qubit Computing Environment).